Given X has a normal distribution with sigma = 50 and standard deviation= 4 then find P(51<=X<=53)=
Answer #1

Solution :

Given that ,

mean = = 50

standard deviation = = 4

P(51 < x < 53) = P[(51 - 50)/ 4) < (x - ) /  < (53 - 50) / 4) ]

= P(0.25 < z < 0.75)

= P(z < 0.75) - P(z < 0.25)

= 0.7734 - 0.5987

= 0.1747

P(51 < x < 53) = 0.1747